OpenStudy (anonymous):

HELP?(: 6 = -2(10n + 7)

OpenStudy (anonymous):

i got n= -1 but I'm not sure that's correct

OpenStudy (anonymous):

Your right

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

6 = -2(10n + 7) 6 = -20n - 14 6 + 14 = -20n 20 = -20n 20/-20 = n -1 = n correct....you got it